4 4. Given the function f(x)- + 3)2 (a) Find any vertical and horizontal asymptotes of the function f(x) (b) Sketch the graph of f(x) 10 10 (c) Find a such that m f(x) 00. a (d) lim /(r) irm
Answer #1

Solution : ( a )

f(x) = -4 (r+3

\mathrm{Vertical\:asymptotes\:of\:}\frac{2}{\left(x+3\right)^2}-4:

\mathrm{Go\:over\:every\:undefined\:point\:}x=a\mathrm{\:and\:check\:if\:at\:least\:one\:of\:the\:following\:statements\:is\:true:}

\lim _{x\to a^-}f\left(x\right)=\pm \infty

\lim _{x\to a^+}f\left(x\right)=\pm \infty

\mathrm{Find\:undefined\:\left(singularity\right)\:points}:

\mathrm{Take\:the\:denominator\left(s\right)\:of\:}\frac{2}{\left(x+3\right)^2}-4\mathrm{\:and\:compare\:to\:zero}

\mathrm{Solve\:}\:\left(x+3\right)^2=0:

\left(x+3\right)^2=0

\mathrm{Using\:the\:Zero\:Factor\:Principle:}

x+3=0

\mathrm{Subtract\:}3\mathrm{\:from\:both\:sides}

x+3-3=0-3

\mathrm{Simplify}

x=-3

\mathrm{The\:final\:solution\:to\:the\:quadratic\:equation\:is:}

x=-3

\mathrm{The\:following\:points\:are\:undefined}

x=-3

\mathrm{The\:vertical\:asymptotes\:are\quad\:}x=-3

\mathrm{and}

\mathrm{Horizontal\:Asymptotes\:of\:}\frac{2}{\left(x+3\right)^2}-4:

\mathrm{Check\:if\:at\:}x\to \pm \infty \mathrm{\:the\:function\:}y=\frac{2}{\left(x+3\right)^2}-4

\mathrm{\:behaves\:as\:a\:line,\:}y=mx+b

\mathrm{Find\:an\:asymptote\:for\:}x\to -\infty \::

\mathrm{Compute}\:\lim _{x\to -\infty \:}\frac{f\left(x\right)}{x}\:\mathrm{to\:find\:m:}

\lim _{x\to -\infty \:}\frac{f\left(x\right)}{x}=\lim _{x\to -\infty \:}\frac{\frac{2}{\left(x+3\right)^2}-4}{x}\quad :

\lim _{x\to \:-\infty \:}\left(\frac{\frac{2}{\left(x+3\right)^2}-4}{x}\right)=0

\mathrm{Now,}

\mathrm{Compute}\:\lim _{x\to -\infty \:}f\left(x\right)-mx\:\mathrm{to\:find\:b:}

\lim _{x\to -\infty \:}f\left(x\right)-mx=\lim _{x\to -\infty \:}\frac{2}{\left(x+3\right)^2}-4-0x\quad :

\lim _{x\to \:-\infty \:}\left(\frac{2}{\left(x+3\right)^2}-4-0*x\right)=-4

y=-4

\mathrm{So,\quad\:}\mathrm{\:an\:asymptote\:for\:}x\to -\infty\mathrm{\: \:is\:}\quad y=-4

\mathrm{and}

\mathrm{Find\:an\:asymptote\:for\:}x\to \infty \::\quad

\mathrm{Compute}\:\lim _{x\to \infty \:}\frac{f\left(x\right)}{x}\:\mathrm{to\:find\:m:}

\lim _{x\to \infty \:}\frac{f\left(x\right)}{x}=\lim _{x\to \infty \:}\frac{\frac{2}{\left(x+3\right)^2}-4}{x}\quad :

\lim _{x\to \infty \:}\left(\frac{\frac{2}{\left(x+3\right)^2}-4}{x}\right)=0

\mathrm{Now,}

\mathrm{Compute}\:\lim _{x\to \infty \:}f\left(x\right)-mx\:\mathrm{to\:find\:b:}

\lim _{x\to \infty \:}f\left(x\right)-mx=\lim _{x\to \infty \:}\frac{2}{\left(x+3\right)^2}-4-0x\quad :

\lim _{x\to \infty \:}\left(\frac{2}{\left(x+3\right)^2}-4-0*x\right)=-4

y=-4

\mathrm{So,\quad\:}\mathrm{\:an\:asymptote\:for\:}x\to \infty\mathrm{\: \:is\:}\quad y=-4

\mathrm{The\:horizontal\:asymptote\:is\quad\:}y=-4

\mathrm{Hence,}

\mathrm{Vertical\:asymptotes\:of\:the\:function\:of\:}f(x):\quad\:x=-3,

\mathrm{Horizontal\:asymptotes\:of\:the\:function\:of\:}f(x):\quad\:y=-4

-----------------------------------------------------------------------------------------------------------------------------

Solution : ( d )

\lim _{x\to \infty \:}f(x)=\lim _{x\to \infty \:}\left(\frac{2}{\left(x+3\right)^2}-4\right)

Steps:

\lim _{x\to \infty \:}\left(\frac{2}{\left(x+3\right)^2}-4\right)

\lim _{x\to a}\left[f\left(x\right)\pm g\left(x\right)\right]=\lim _{x\to a}f\left(x\right)\pm \lim _{x\to a}g\left(x\right)

With the exception of indeterminate form

=\lim _{x\to \infty \:}\left(\frac{2}{\left(x+3\right)^2}\right)-\lim _{x\to \infty \:}\left(4\right)

=2* \lim \:_{x\to \infty \:}\left(\frac{1}{\left(x+3\right)^2}\right)-\lim _{x\to \infty \:}\left(4\right)

=2* \frac{\lim _{x\to \infty \:}\left(1\right)}{\lim _{x\to \infty \:}\left(\left(x+3\right)^2\right)}-\lim _{x\to \infty \:}\left(4\right)

=2* \frac{1}{\infty^{2} \:}-4

=2* \frac{1}{\infty \:}-4

=2* 0-4

=0-4

=-4

\mathrm{Hence,}

\lim _{x\to \infty \:}f(x)=-4

-------------------------------------------------------------------------------------------------------------------

Solution : ( e )

\lim _{x\to -\infty \:}f(x)=\lim _{x\to -\infty \:}\left(\frac{2}{\left(x+3\right)^2}-4\right)

Steps:

\lim _{x\to -\infty \:}\left(\frac{2}{\left(x+3\right)^2}-4\right)

=\lim _{x\to \:-\infty \:}\left(\frac{2}{x^2+6x+9}-4\right)

\lim _{x\to a}\left[f\left(x\right)\pm g\left(x\right)\right]=\lim _{x\to a}f\left(x\right)\pm \lim _{x\to a}g\left(x\right)

With the exception of indeterminate form

=\lim _{x\to \:-\infty \:}\left(\frac{2}{x^2+6x+9}\right)-\lim _{x\to \:-\infty \:}\left(4\right)

=2*\lim \:_{x\to \:-\infty \:}\left(\frac{1}{x^2+6x+9}\right)-\lim _{x\to \:-\infty \:}\left(4\right)

=2*\frac{\lim _{x\to \:-\infty \:}\left(1\right)}{\lim _{x\to \:-\infty \:}\left(x^2+6x+9\right)}-\lim _{x\to \:-\infty \:}\left(4\right)

=2*\frac{1}{\infty \:}-4

=2*0-4

=0-4

=-4

\mathrm{Hence,}

\lim _{x\to -\infty \:}f(x)=-4